# Doc-PP: Document Policy Preservation Benchmark for Large Vision-Language Models

### v1

@@ Abstract
The deployment of Large Vision-Language Models (LVLMs) for real-world document question answering is often constrained by dynamic, user-defined policies that dictate information disclosure based on context. While ensuring adherence to these explicit constraints is critical, existing safety research primarily focuses on implicit social norms or text-only settings, overlooking the complexities of multimodal documents. In this paper, we introduce Doc-PP (Document Policy Preservation Benchmark), a novel benchmark constructed from real-world reports requiring reasoning across heterogeneous visual and textual elements under strict non-disclosure policies. Our evaluation highlights a systemic Reasoning-Induced Safety Gap: models frequently leak sensitive information when answers must be inferred through complex synthesis or aggregated across modalities, effectively circumventing existing safety constraints. Furthermore, we identify that providing extracted text improves perception but inadvertently facilitates leakage. To address these vulnerabilities, we propose DVA (Decompose–Verify–Aggregation), a structural inference framework that decouples reasoning from policy verification. Experimental results demonstrate that DVA significantly outperforms standard prompting defenses, offering a robust baseline for policy-compliant document understanding. [^1] https://github.com/hwanchang00/doc-pp
@@
Haeun Jangthanks:   Equal contribution.   Hwan Chang11footnotemark: 1   Hwanhee Leethanks:   Corresponding author. Chung-Ang University, Seoul, Korea {jhe020814,hwanchang,hwanheelee}@cau.ac.kr
@@ 2 Related Work
As large language models (LLMs) (OpenAI, 2025b) have advanced, they are increasingly deployed in settings where rich context is provided at inference time, often including personal, organizational, or situational information. This has motivated a growing body of work on how models handle private or sensitive information conditioned on the surrounding context. Prior studies such as ConfAIde (Mireshghallah et al., 2024)and PrivacyLens (Shao et al., 2024) investigate whether LLMs respect contextual privacy expectations, drawing on the theory of Contextual Integrity (Nissenbaum, 2004) to evaluate the appropriateness of information flow under different social roles, recipients, and interaction settings. Recently, this line of inquiry has expanded to the multimodal domain. VLM-GEOPRIVACY Anonymous (2025) investigates whether LVLMs can respect privacy in geolocation tasks by interpreting visual cues. However, these works predominantly focus on implicit norms of information disclosure—relying on the model to deduce what is appropriate based on broad social conventions or assumed privacy expectations—rather than enforcing specific directives provided by the user. In parallel, CoPriva Chang et al. (2025) shifts the focus to explicit, user-defined constraints, benchmarking model adherence to specified non-disclosure policies in text-only environments. Our work extends this frontier into the multimodal document domain. Unlike prior text-based benchmarks, Doc-PP requires models to perform cross-modal integration to identify and prevent policy violations, particularly in scenarios where sensitive information is dispersed across complex visual-linguistic dependencies.
@@ 3.2 Construction Pipeline
The overall construction pipeline of Doc-PP Bench follows three-stage process—comprising policy construction, query construction, and evaluation—as illustrated in Figure  2. We provide full prompts in Appendix F.
@@ Results.
On Gemini-3-Flash-Preview, DVA substantially reduces leakage on our benchmark, lowering the leakage rate from 64.6 (default) to 30.5. This result suggests that the bottleneck in multimodal policy preservation is the granularity of verification; by isolating claims, DVA transforms the task of evaluating long-form responses into simpler, atomic verification tasks, thereby preventing leaks that are otherwise overlooked in integrated outputs. As demonstrated in the case study in Figure 4, DVA’s element-level verification prevents factual leaks that remain undetected in both CoT and revision-based baselines. Additional failure cases are detailed in the Appendix E.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)] and the National Research Foundation of Korea(NRF) grantfunded by the Korea government(MSIT) (RS-2025-24683575).
@@ Appendix D Additional Results
We examine whether increasing the context size obscures sensitive information, potentially reducing leakage. Table 6 shows the leakage rates across varying context lengths, from 2 to 11 surrounding pages. Counterintuitively, increasing the context window does not provide a safety buffer; leakage rates remain consistently high (fluctuating between 60% and 70%) regardless of document length. This suggests that leakage is not primarily driven by long-range context accumulation, but rather by localized cross-modal reasoning within a small subset of relevant pages.
@@
Appendix E Failure Cases Analysis
@@
Appendix F Prompt Templates
@@
F.1 Policy Generation Prompt
@@
F.2 Value Extraction Prompt
@@
F.3 Restricted Target Checklist Verifier Prompt
@@
F.4 Implicit Query Generation Prompt
@@
F.5 Content-Coverage Checklist Generation Prompt
@@
F.6 Content-Coverage Checklist Validation Prompt
@@ F.7 Evaluation Prompt
F.8 Implicit Leakage and Evaluation Prompt
@@
F.9 Chain-of-Thought (CoT) Prompt
@@
F.10 Post-Hoc Revision Prompt
@@
F.11 DVA (Decompose–Verify–Aggregation) Framework Prompt

### v2

@@ Abstract
The deployment of Large Vision-Language Models (LVLMs) for real-world document question answering is often constrained by dynamic, user-defined policies that dictate information disclosure based on context. While ensuring adherence to these explicit constraints is critical, existing safety research primarily focuses on implicit social norms or text-only settings, overlooking the complexities of multimodal documents. In this paper, we introduce Doc-PP (Document Policy Preservation Benchmark), a novel benchmark constructed from real-world reports requiring reasoning across heterogeneous visual and textual elements under strict non-disclosure policies. Our evaluation highlights a systemic Reasoning-Induced Safety Gap: models frequently leak sensitive information when answers must be inferred through complex synthesis or aggregated across modalities, effectively circumventing existing safety constraints. Furthermore, we identify that providing extracted text improves perception but inadvertently facilitates leakage. To address these vulnerabilities, we propose DVA (Decompose–Verify–Aggregation), a structural inference framework that decouples reasoning from policy verification. Experimental results demonstrate that DVA significantly outperforms standard prompting defenses, offering a robust baseline for policy-compliant document understanding.
@@
Haeun Jangthanks:   Equal contribution.   Hwan Chang11footnotemark: 1   Hwanhee Leethanks:   Corresponding author. Chung-Ang University, Seoul, Korea {jhe020814,hwanchang,hwanheelee}@cau.ac.kr https://hwanchang00.github.io/docpp_project_page
@@ 2 Related Work
As large language models (LLMs) (OpenAI, 2025b) have advanced, they are increasingly deployed in settings where rich context is provided at inference time, often including personal, organizational, or situational information. This has motivated a growing body of work on how models handle private or sensitive information conditioned on the surrounding context. Prior studies such as ConfAIde (Mireshghallah et al., 2024)and PrivacyLens (Shao et al., 2024) investigate whether LLMs respect contextual privacy expectations, drawing on the theory of Contextual Integrity (Nissenbaum, 2004) to evaluate the appropriateness of information flow under different social roles, recipients, and interaction settings. Recently, this line of inquiry has expanded to the multimodal domain. VLM-GEOPRIVACY Anonymous (2025) investigates whether LVLMs can respect privacy in geolocation tasks by interpreting visual cues. However, these works predominantly focus on implicit norms of information disclosure—relying on the model to deduce what is appropriate based on broad social conventions or assumed privacy expectations—rather than enforcing specific directives provided by the user. In parallel, CoPriva Chang et al. (2025) shifts the focus to explicit, user-defined constraints, benchmarking model adherence to specified non-disclosure policies in text-only environments. Our work extends this frontier into the multimodal document domain. Unlike prior text-based benchmarks, Doc-PP requires models to perform cross-modal integration to identify and prevent policy violations, particularly in scenarios where sensitive information is dispersed across complex visual-linguistic dependencies. A detailed feature-level comparison between CoPriva and Doc-PP is provided in Appendix E.
@@ 3.2 Construction Pipeline
The overall construction pipeline of Doc-PP Bench follows three-stage process—comprising policy construction, query construction, and evaluation—as illustrated in Figure  2. We provide full prompts in Appendix H.
@@ Results.
On Gemini-3-Flash-Preview, DVA substantially reduces leakage on our benchmark, lowering the leakage rate from 64.6 (default) to 30.5. This result suggests that the bottleneck in multimodal policy preservation is the granularity of verification; by isolating claims, DVA transforms the task of evaluating long-form responses into simpler, atomic verification tasks, thereby preventing leaks that are otherwise overlooked in integrated outputs. As demonstrated in the case study in Figure 4, DVA’s element-level verification prevents factual leaks that remain undetected in both CoT and revision-based baselines. Additional failure cases are detailed in the Appendix G. A further analysis of DVA’s faithfulness scores and response verbosity is provided in Appendix F.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)] and the National Research Foundation of Korea(NRF) grantfunded by the Korea government(MSIT) (RS-2026-25494299).
@@ Appendix D Additional Results
We examine whether increasing the context size obscures sensitive information, potentially reducing leakage. Table 6 shows the leakage rates across varying context lengths, from 2 to 25 surrounding pages. Counterintuitively, increasing the context window does not provide a safety buffer; leakage rates remain consistently high (fluctuating between 60% and 70%) regardless of document length. This suggests that leakage is not primarily driven by long-range context accumulation, but rather by localized cross-modal reasoning within a small subset of relevant pages.
@@
Appendix E Comparison with CoPriva
Table 7 summarizes the key differences between CoPriva (Chang et al., 2025), a text-only policy QA benchmark, and our Doc-PP benchmark, which operates in the multimodal document domain.
Feature CoPriva (Text-only Policy QA) Doc-PP (Multimodal Policy QA)
Evidence Scope Localized to short, contiguous meeting transcript spans. Multi-span / Multi-page synthesis: Requires tracking and aggregating fragmented evidence across extensive documents.
Nature of Leakage Direct or semantic extraction of a pre-existing discussion summary. Compositional inference: Models must dynamically compute or deduce hidden values by combining heterogeneous modalities.
Target Granularity Abstract/semantic topics. Value-centric targets: Grounded in specific, exact values that require computational reasoning to uncover.
Query Construction Adapted: Uses existing summary queries to test implicit leakage. Value-dependent generation: Generates novel implicit queries strictly designed so that a faithful summary computationally forces disclosure of the protected exact value.
Evaluation Metric LLM-as-a-judge checking for semantic inclusion of the policy topic. Strict value verification: Leakage is strictly defined by the presence of the computed value $v$, coupled with a rigorous content-coverage checklist.
Table 7: Feature-level comparison between CoPriva and Doc-$P$.
Appendix F DVA Faithfulness and Verbosity Analysis
We further analyze the trade-offs of each mitigation strategy in terms of faithfulness and response verbosity. Table 8 reports the average number of tokens generated by each method, and Table 9 presents the leakage and faithfulness scores.
Method Avg. Tokens
Naive 382.42
CoT 551.67
Revision 862.36
DVA 935.97
Table 8: Average response length (in tokens) for each mitigation method.
Method Leakage ($\downarrow$) Faithfulness ($\uparrow$)
CoT 51.0 59.0
Revision 38.8 66.9
DVA 30.5 55.8
Table 9: Leakage and faithfulness scores for each mitigation method.
Appendix G Failure Cases Analysis
@@
Appendix H Prompt Templates
@@
H.1 Policy Generation Prompt
@@
H.2 Value Extraction Prompt
@@
H.3 Restricted Target Checklist Verifier Prompt
@@
H.4 Implicit Query Generation Prompt
@@
H.5 Content-Coverage Checklist Generation Prompt
@@
H.6 Content-Coverage Checklist Validation Prompt
@@ H.7 Evaluation Prompt
H.8 Implicit Leakage and Evaluation Prompt
@@
H.9 Chain-of-Thought (CoT) Prompt
@@
H.10 Post-Hoc Revision Prompt
@@
H.11 DVA (Decompose–Verify–Aggregation) Framework Prompt